RG004-AsH  Gas Detectors resistive sensors

RG004 is a family of thin-film resistive gas sensors tailored for detection of small quantities of specific substances dissolved into air. The thin-film layer is activated with an organic molecule that guarantee high sensibility, high specificity and relative immunity from temperature and humidity variations. RG004-AsH is designed to detect average concentration of arsine; suitable to monitot gas escapes for semiconductor growth facilities, MBE and MO-CVD machines, etc.

Electrical Characteristics

a) sens. slope @ 0 ppm

b) suggested lineal approx.

c) real calibration curve

R(C) = R_{0}e^{−C/C_{0}}

Min Typ Max Units
Max range  100 ppm
Resistance @ 0 ppm  10  kΩ
Sensitivity @ 0 ppm  -40  Ω/ppm

R_{0} accuracy (full temp range)

1 %

Step response time (to 90%)

90 msec

Temperature range  -55–240  ^{∘}C

Maximum power

dissipation

80 mW

Packaging

Notice: 3-4 is a temperature-drift matched resistor of nominal R_{0}, useful for high-sensitivity low range sensing.